UT 03 Jun 2020

Restart Night Log for C19 Restart

OSA: D. Gonzalez Huerta (LBTO Remote Room), J. Hill (Tucson)

ISA: O. Kuhn (Tucson)

Nighttime Mtn Assistant: T. Forsyth

Mtn Manager: R. Hansen


Night Plan

  • Telescope Check
    • Test preset telescope
    • Open/Close Shutter Doors
    • Test Polycom w/Zoom setup
  • LBC to start if sunset is clear - Skyflats
  • MODS to start if cloudy at sunset
  • MODS AGW checkouts
  • MODS test OBs
    • Q1704+710
    • Photometric Standards
    • Spectrophotometric Standards
  • LBC collimation
  • LBC superfoc
  • LBC test OBs (sidereal and NS)
    • MCG+05-40-000 17:12:12.0 28:25:00.0 - up all night
      • SDSS and Bessel
    • Dumbell Nebula (11pm onwards)
    • C/2017 T2 (Panstarrs) NS - start of night
    • C/2019 K7 (Smith) NS - end of night
    • Photometric Standards
    • Spectrophotometric Standards
  • LBC ITs

Summary

Last night we completed a basic checkout of MODS and the LBCs. We started with MODS and observed a spectrophotometric standard in dual + red grating and dual prism modes, we made a long-slit science observation and then a non-sidereal dithered & guided imaging observation. We ran into a copointing check violation during the non-sidereal observation, which was resolved by disabling the copointing check. We switched to the LBCs after midnight, but still started with the glass about 3 deg warmer than the ambient. The difference was down to 2 deg the last time I checked, but still collimation (dohybrid and dofpia) would not converge and we could see the higher order spherical building up. I stopped each run of FPIA with a shift-S just so that we could move on to take some science-like data. The seeing was ~1" or less, but IQ was poor and changing. We took some guided & dithered sidereal images at Bessell & SDSS filters, then made a non-sidereal guided & dithered observation. Finally we ended with skyflats. We obtained some good flats at each of the filters: in LBCB, g, V, B and in LBCR, r and R. The night did not start out photometric, but I think it was clear during the second half. I did not take LBC standards, but I think that photometry from the images might be good. I didn't see clouds at sunrise or on the allsky during the 2nd half.

Some problems aside from the temperature-related problems with LBC collimation:
  • mods1r is showing it seems a quite high frequency of bad reads: 5 in the past 24 hours. Notes added to IT 5862.
  • an error with the copointing check for the MODS NS observation
  • No IT or emails sent yet about the following:
    • telescope logs not where LBTPlot expects them (there is no 2020/06 directory under /lbt/logs) so the LBC guiding corrections do not show up in LBTplot.
    • permissions issue preventing viewing Tucson Repository for all dates since restart (can access 20200318 and earlier, but not 20200529 and later)
    • a screenshot caused rm507-1 to hang

On the whole, it was a nice night and the communications and connections all worked pretty well, from my perspective. I am not sure how others perceived it, though.

Log

01:30 MODS1 Blue fiber swapped, seqinit. MODS both awake and sieveSnaps being taken.

01:30 Turning on the LBCs.

01:50 Sky does not look so bad. I am going to take MODS twilight flats with a comment in the header. Running a mods script to see how long the PROPID can be. "abcdefghijk-2020A-lbt-mods/LBT-2019B-008" does not crash the system. There is an LBTB script that has a very long PROPID.

02:27 Terry, John H, David GH and I (plus Josh as a spectator) are all on the polycom & zoom: lbtopolycom. Connection! (correction... we did not yet have the rm507 polycom connection. We are bringing that in now, 03:55 UT)

02:34 clouds around us, but above it is pretty nice. Error on a vent door, but that is not delaying opening. Delaying opening for some clouds.

03:05 Opening!!

03:15 execBinoMODS gi_Dual_1Ktest_Blank16+55.txt ... just a test of the pair of SkyFlat scripts: 1K test (gen from OT) and 3K observing (from my template in ~modseng/modsScripts/modsCalib/SkyFlats/ SkyFlat scripts.
  • execBinoMODS gi_Dual_1Ktest_Blank16+55.txt
  • execMODS --mods1 --bino -e gi_skyflat_obs.txt
    • "-e" was used to skip to exec block and the test was here to insure that not filling in the coords, but leaving them as hh:mm:ss dd:mm:ss did not matter.
  • test also checks that no guide filter, no guide star coords does not cause a problem for the track preset. It does not.

03:29 Slewing to a pointing star near HZ44. There is cirrus, but we'll observe this in dual, red grating and dual prism modes.

03:35 David is reporting some GCSGUI sluggishness. John, David and I are all running GCSGUIs through x2go's. I see both GCSGUIs (x2go on to obs3) and they are updating. John and I closed our GUIs in the hope that David's GUIs would speed up and they seem to be working better on his setup now.

03:49 Closing for clouds.

04:32 Reopening. On the mountain, they lost the HBS, polycom, allsky + weather for a while.

04:49 Back to a pointing star near HZ44.

04:58 Collimation off-axis. HBS oil is heating up. TA 22 C, but ambient temp is ~12 C. Dan is assisting.

05:22 acqBinoMODS hz44acq.txt
  • mods1r.20200603.0009.fits shows a readout error. 2nd one today.
  • modsAlign -r mods1r.20200603.0010.fits
    • MODS1 Offset Command: offsetxy -0.206 11.170 rel, tweaking by offsetxy -0.378 0.221 rel
  • modsAlign -r mods2r.20200603.0010.fits
    • MODS2 Offset Command: offsetxy 4.506 7.295 rel, tweaking by offsetxy -0.424 0.256 rel
  • 0012 will be the 2nd set of confirmatory thru-slit images.
    • better.

05:37 execBinoMODS hz44dualgrating.txt, followed by hz44redgrating.txt

06:00 bringing up my GCSGUIs again. Seeing is ~0.8".

06:22 acqBinoMODS q1704acq1.txt PA=22 deg. Guide star not found on right side and so David will correct the pointing here.

06:31 Retrying after pointing correction.
  • modsAlign 1 offsetxy -0.506 10.851 rel
  • modsAlign 2 offsetxy 4.343 7.042 rel
  • gave -2 pix adjustment to mods1 and -0.15" to mods2 --> 25
06:45 Taking a series of 3 5-min exposures on the target. Seeing is about 1.1" here.
  • mods1r: 26-28, mods1b: 15-17
  • mods2r:

07:11 NS test. Slewing to 2017 UR4 to take a series of guided, dithered MODS images.
  • after first exposure at 10, 10, it will pause and I will offset more
  • blocked; try to shorted exposure time. did so
  • mods2b 18, mods2r 28, mods1b 18, mods1r 29 (3rd seq problem today). repeating mods1r
  • tried a 10" 10" detxy offset and then updated pointing
  • resume dither sequence, so now it is going to -10, -10 but exposing. MODS is in a race condition, aborting.

07:37: Trying again, but this time with the lines to make a syncoffset to the upper left quadrant and then updatepointing before starting the series of exposures.
  • set WFS exptime to 7000ms
  • m1b 19-21, m1r 31, m2b 20, m2r 30.
  • 07:45 UT. copointing limit check violated on -10 -10 offset (before m2b22 and m2r32), but we wonder why. The script made a -20 40 sync offset, then issued an updatepointng, after which it was commanded to dither by (detxy abs): (-10", 10"), (-10", -10"), (+10", -10") (see below). David diabled the copointing limit check and I retried the offset. It worked.
  syncoffset 
  offsetxy -20 40 abs
  updatepointing

  #OFFSETXY  10.000 10.000 abs
  BLUE FILTER g_sdss
  BLUE EXPTIME 60.0
  BLUE NIMGS 1
  BLUE CCDBIN 1 1
  RED FILTER r_sdss
  RED EXPTIME 60.0
  RED NIMGS 1
  RED CCDBIN 1 1
  WFSwait
  GO
  OFFSETXY  -10.000 10.000 abs
  GO
  OFFSETXY  -10.000 -10.000 abs   <---  co-pointingLimitError performed. Target on side right is greater than 51" from the computed midpoint: Error limit is 50.
  GO
  OFFSETXY  10.000 -10.000 abs
  GO

07:55: Reconfiguring for LBCs

08:22 Sending preset to the copointing field for our LBC field for the night, MCG+05-40-000 (17:11, +28d)

08:30 dohybrid.
Waiting for file in /newdata/lbc?.20200603.??????.fits
******* WARNING *******
Data files DID NOT appear in /newdata
Pls double check and then: press Q to exit or any other key to continue

First images had strong coma. b John applied -0.8mm X on both sides, but I then allowed WRS to analyze these images and send the corrections, so the results will not be very meaningful.

08:38 Clearing active optics. John applying -0.8mm X on both sides. Now I am running dohybrid (or perhaps I just ran dofpia)
  • b 083946 and r 083941. These appear to have less coma than before, but they still have coma.
  • Collimation is not converging, we are seeing high order spherical. The mirrors are 3C above the ambient. I'm stopping this with "S" but I was too late.

08:53 Clearing active optics again. The -0.8mm X global offset is in. John is turning off the Z22 gains.
  • dohybrid. The first blue image has z7 now, but the DX side has z8. WRS introduced astigmatism (see 085752). We may need to turn off astig in WRS.
  • dohybrid thinks it converged, but blue definitely has not. It still has astig , 090036.

09:00 Blue said it converged but it had not - it found only 1 pupil and then rejected it.

09:00 We are clearing active optics and repeating dohybrid. Left: glass 15, ambient 12.8. This time it looks like wrs did not introduce astig.
  • Blue side - all pupils are being rejected. Could be because the background, with the moon, is too high. Giving a "S" to stop gracefully.

09:14 Launching a science-like OB which will guide and dither.
  • Us,g,r and z,i,r
  • 091442 and 091432
  • guiding is working, although LBTplot is not showing it because the logs are not on obs1 in the expected location (they are on tcs1, just not on obsN).
  • dithering is working. compare 092402 with 091432.

09:35 Clearing active optics. Collimating again. I'd forgotten to copoint so will do that and take a series at Bessell filters.
  • 094010, 094002 are the first fpia...
  • "S" to stop it because it is not converging.
  • CP images: 094327 and 094323
    • offsets SX: dX = -0.255, dY = -0.06, dRX = -1.287, dRY = 5.5
    • offsets DX: dX = +0.200, dY = -0.111, dRX = -2.384, dRY = -4.282
09:49 Run a few Bessel exposures.

10:20 A printscreen issued on rm507-1 hung the computer. Trying to find a process to kill or recover the x2go session. David recovered this by terminating old x2go and starting a new one - he found everything up and running.

10:25 Sending preset to the copointing star for 2017 UR4.
  • dofpia --- we never really achieved collimation, but temps are still out of whack with glass ~ 2 deg warmer than ambient.
  • a VERY crowded field for collimation. The first pupils do not look so bad, but z22 corrections are -100, so I am "S"ing to stop it gracefully.

10:32 Sending the NS OB (edited to make PA = 0 for both sides). This will take a series of 9 dithered 60-sec exposures on both sides.

10:37 Stopping to try to collimate better. It seems that it was going in the right direction earlier.
  • It was again giving -ve Z22 and not converging. I stopped it again and will run the OB.

10:44 Copointing. b 104634 r 104625

10:50 Running the science OB on 2017 UR4. Starting with b 105145 and r 105137. Sidereal. Red IQ is poor. Blue is looking OK.

10:56. Running the science OB but NS.

10:58 Trying to collimate to improve Red, but it is not going, really.

11:00 One more try on the NS OB, this time with NS tracking.

11:17 Rotational trailing test.
  • Slew to copoint/collimation field at RA~LST and DEC=29:24.
  • dofpia, ended by "S"
  • copoint ... sky is already 17th mag. Abandoning this to get sky flats

11:29 Slew to Blank 19+55
  • gr test ~8000
  • gr dither 3x then 1x. 3 good ones at g, and 2 (maybe 3) at r
  • VR 0deg 0.5x starting with 113841 and 113833 (11k - 17k and 8k - 15k counts)
  • BR 180deg 0.5x starting with 114332 and 114329 (22k 45 - and 24k - 48)
  • missed Us and z flats, too bright.

11:52 Closing up. Will do some cals. Note, the debris at the bottom left corner of the MODS1 images is gone.
  • LBC 25 Biases, 3x
Type MODS1 B MODS1 R MODS2 B MODS2 R
biases 8K 23-27 35-39 24-28 34-38
biases 4K 28-32 40-44 29-33 39-43
biases 3K 33-37 45-49 34-38 44-48
biases 1K 38-42 50-54 39-43 49-53
imflats dual u & r 43-47 55-59 44-48 54-58
imflats dual g & i 48-52 60-64 49-53 59-64
imflats dual z   65-69   65-69
imflats red r,i,z   70-74,75-79,80-84   70,71, nimgs in mods2 red script is 1 not 5
imflats red r,i,z       72,73,74
imflats red r,i,z       75-77,78-80,81-83
grlamps dual 53-55,56(rep) 85-87,88(rep) 54-56 84-86
prlamps dual 57-60,61(rep) 89-92,93(rep)    
pixflats dual 62-66,67-71 94-99 57-61 87-91
14:40 Done with calibrations.

TBD still:
  • MODS 5" and 1" slit flats dual
  • MODS 5" red
  • MODS pix flats red
  • MODS lamps: red

The TBD cals were done on 20200606.
  • MODS 5" dual slit flats: m1b: 3-8, m1r: 3-5, m2b: 3-8, m2r: 3-5
  • MODS 1" dual slit flats: m1b: 9-14, m1r: 6-11, m2b: 9-14, m2r: 6-11
  • MODS 5" red slit flats: m1r: 12-14 , m2r: 12-14
  • pix flats red: m1r: 15-19, m2r: 15-19
  • grlamps red: m1r: 20-22,24-26, m2r: 20-22
    • m1r 20 - Ne lamp did not fire! sigh... again??. Repeated with a test (23) and the sequence (24), Ne came on for these two additional exposures. I noticed it was on when I checked MODS last night around 22:00, so it could have been on for ~10 hours, since the test through the PCI machine which finished ~12:20pm Friday.

mods1r bad reads: <br 20200602.0005 (on Ne test)
20200603.0009
20200603.0029
20200603.0085 (on Ne+Hg)
20200603.0089 (also on Ne+Hg)

-- OlgaKuhn - 03 Jun 2020
Topic revision: r5 - 06 Jun 2020, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback